OpenStudy (anonymous):

These lines are perpendicular. Is this statement true or false? y =-2/5x + 4 y =5/2x − 3

OpenStudy (loser66):

if I said it's true, do you trust me? is there any way to test?

OpenStudy (anonymous):

sure and idk

OpenStudy (loser66):

I told you before, if 2 lines perpendicular, then the slope is backward let say m1 is the slope of line 1 m2 is the slope of line 2 then m2= -1/m1

OpenStudy (loser66):

now, check, m1 = -2/5 m2 = -1/m1 = \(\dfrac{-1}{-2/5} = \dfrac{5}{2}\) so that, they are?

OpenStudy (anonymous):

srry i told you im an idiot and whats that mean

OpenStudy (loser66):

That's clear as you know your own name. \(m_2 =\dfrac{-1}{m_1}\) so, just plug those m's in \(\dfrac{5}{2}= \dfrac{-1}{-2/5}\) and it works

OpenStudy (anonymous):

k